# MultiFinBen: A Multilingual, Multimodal, and Difficulty-Aware Benchmark for Financial LLM Evaluation

### v2

@@ F.1 English
For the IE task, it focuses on extracting structured financial information from unstructured text. We evaluate this capability using three datasets: SC [43], FinRED [62], and FINER-ORD [61]. The SC [43] dataset assesses the detection of causal relationships in financial news and SEC filings, using F1 [65] as the evaluation metric. The FinRED [62] dataset targets relation extraction, identifying relationships such as “product/material produced” and “manufacturer” in financial news and earnings call transcripts, also evaluated using F1 [65]. The FINER-ORD [61] dataset involves named entity recognition, extracting entities like organizations, locations, and persons from financial agreements and SEC filings, with performance measured by entity F1 [12]. Overall, the IE task provides a foundation for evaluating the model’s understanding of financial knowledge.
@@
For the TA task, it encompasses tasks such as sentiment classification, topic detection, and event identification from financial texts. In this work, we focus on two representative financial datasets: Headlines [64] and TSA [11]. The Headlines [64] dataset targets the extraction of actionable signals—such as price movements—from financial news headlines, with performance measured by the average F1 score (avg F1). The TSA [11] dataset centers on sentiment analysis, where the goal is to classify textual segments into positive, negative, or neutral sentiment categories, evaluated using the accuracy [42] metric. This task is designed to assess the capability of LLMs in performing fine-grained understanding and reasoning over domain-specific financial language.
@@
For the QA task, it involves answering specific regulatory and financial questions by leveraging the inherent knowledge of LLMs. Both FinQA [8] and TATQA [92] assess models’ numerical reasoning abilities using financial reports, tables, and contexts, whereas XBRL-Math [9] focus on equation inference. All are evaluated by accuracy [42].
@@
For the TG task, it involves summarizing or explaining specific content within regulatory or financial contexts, evaluating models’ ability to produce coherent and informative text. EDTSUM [79] targets the summarization of financial news articles into concise and informative summaries. In addition, we utilize the ECTSUM [45] dataset, which focuses on summarizing earnings call transcripts. Model performance on these summarization tasks is evaluated using ROUGE-1 [37].
@@
For the RM task, it focuses on identifying, extracting, and analyzing risk-related information, interpreting numerical data, and evaluating a model’s ability to understand complex relationships. Specifically, we use the CCF [14] datasets to assess LLMs’ ability in our paper. The CCF [14] dataset is suitable for detecting fraud, which identifies whether transactions are “fraudulent” or “non-fraudulent”. This dataset is evaluated by Matthews Correlation Coefficient (MCC) [44].
@@
For the FO task, it evaluates a model’s ability to predict future market trends and investor behavior based on patterns in financial data. We focus on stock movement prediction, aiming to forecast whether a stock’s price will rise or fall using historical price data and social media signals from the BigData22 [66] dataset. Model performance is assessed using MCC [44].
@@
For the DM task, it evaluates an LLM’s ability to synthesize diverse financial information to formulate and execute trading strategies—an inherently challenging task even for human experts. We refer to MSFT [85] dataset, which consist of data over one year, simulating real-world trading scenarios using historical prices, news articles, and sentiment analysis. Model performance is assessed using Sharpe Ratio (SR) [63], providing a comprehensive evaluation of profitability, risk management, and strategic decision-making.
F.2 Chinese
@@
In IE task, RRE [23] is a relation extraction task that identifies the semantic relation between two audit entities in a sentence. Relations span 8 predefined categories including audit issue, audit item, audit basis, audit method, audit institution, audit outcome, audited entity, and related sector or domain. This task tests a model’s ability to infer structured relationships in financial text.
@@
For TA task, AIE [23] involves categorizing audit-target entities into 7 regulatory domains (fiscal audit, infrastructure audit, customs audit, financial audit, social security audit, tax audit, and environmental audit). This task evaluates regulatory disambiguation within Chinese financial texts. LNE [23] is a legal grounding task where models must align a given audit scenario with the most relevant legal or regulatory category (financial regulations, fiscal law, personal income tax, financial supervision, labor and employment, listed companies, social security, industry regulation, value-added tax, tax administration, asset evaluation law, resource tax, integrated governance, and comprehensive tax policy). It requires both regulatory reasoning and knowledge of Chinese legal terminology.
@@
In QA task, FinanceIQ [89] is a multi-choice classification task where models need to answer Chinese-language financial domain-specific questions. The task evaluates an LLM’s understanding of domain-specific vocabulary, foundational financial knowledge, and contextual reasoning in Chinese.
All datasets are evaluated using classification accuracy [42].
F.3 Japanese
@@
To capture the unique challenges posed by Japanese in financial contexts, we only find one public TA dataset from a Japanese Financial Benchmark [21] specifically designed to evaluate Japanese financial language understanding.
@@
For TA task, we adopt the chabsa dataset (chABSA: Aspect-Based Sentiment Analysis dataset in Japanese) [33], a binary classification task evaluates the sentiment polarity (positive or negative) of financial terms within securities reports. It is derived from publicly disclosed annual securities filings in Japan. The dataset includes annotated sentiment terms and is scored using the macro F1 [20] metric, excluding neutral terms for evaluation consistency.
F.4 Spanish
@@
To comprehensively evaluate the understanding capabilities of LLMs in Spanish regulatory and financial contexts, we included four domain-specific datasets from FLARE-ES [88] across three tasks: TA, QA, and TG.
@@
For TA task, text classification and sentiment analysis were included. In text classification task, the MultiFin [30] dataset evaluates the model’s ability to categorize Spanish financial texts. This dataset, which centers on Spanish headlines, contains 2,066 articles spanning six key financial categories: Business & Management, Finance, Government & Controls, Industry, Tax & Accounting, and Technology. The task challenges the model to accurately assign each headline to its respective sector, thereby testing its linguistic adaptability and domain-specific knowledge in Spanish. In sentiment analysis task, the TSA [51] dataset includes 3,892 entries from financial news and tweets, annotated to classify sentiments as positive, negative, or neutral—capturing a nuanced spectrum of market emotions in Spanish.
@@
For QA task, the EFPA dataset, derived from questions used in financial examinations by official examiner associations, present unique challenges. the EFPA dataset broadens this challenge with 228 questions featuring four answer choices each, covering a broader spectrum of financial topics including economic principles, fundamental financial concepts, and intricate calculations associated with financial products. This dataset is evaluated using accuracy [42].
@@
For TG task, the FNS-2023 [86] dataset, consisting of 232 Spanish annual reports from diverse financial companies. It aims to condense voluminous Spanish financial documents into succinct, informative abstracts, enhancing the accessibility and usability of financial information. Rich in detailed financial data and narratives, these reports pose a distinct challenge: distilling the essence of each document into a summary that captures the key information while preserving the factual integrity and coherence of the original text. This dataset is evaluated using ROUGE-1 [37].
F.5 Greek
@@
We adopt four core financial NLP datasets from Plutus-ben [54], the first Greek financial benchmark. These datasets span from IE, TA, QA, and TG tasks.
@@
For IE task, only GRFinNUM [54] is included. GRFinNUM [54] targets fine-grained classification of numerals into MONETARY, PERCENTAGE, TEMPORAL, QUANTITY, and OTHERS categories, which captures the nuances of numerical semantics in long-form Greek financial texts. This dataset is evaluated using entity F1 [12].
@@
For TA task, GRMultiFin [31] focuses on classifying concise Greek financial headlines into one of six predefined thematic categories: Business & Management, Finance, Government & Controls, Tax & Accounting, Technology, and Industry, emphasizing the need for lexical disambiguation and contextual inference due to the brevity and ambiguity typical of headlines. This dataset is evaluated using accuracy [42].
@@
For QA task, GRFinQA [54] evaluates models’ performance on inferring the correct answer using provided financial text under a multiple-choice format, assessing a model’s ability to comprehend and reason over finance-related questions drawn from Greek academic sources. This dataset is evaluated using accuracy [42].
@@
For TG task, GRFNS-2023 [86] tests models’ ability of generating fluent, coherent, and factually accurate summaries from complex narrative sections of Greek financial reports, testing a model’s capacity for paraphrasing and information compression. This dataset is evaluated using ROUGE-1 [37].
Appendix G Audio Benchmark
@@
The Automatic Speech Recognition (ASR) task evaluates AudioLLMs’ accuracy in transcribing financial audio clips, directly impacting applications such as financial voice assistants. In this task, each input is represented as $(A,Q,R)$, where $A$ is an audio clip, $Q$ is the prompt instruction, and $R$ is the reference transcript. The transcribed text $T$ is generated as $T=\textit{AudioLLM}(A,Q)$. The Word Error Rate (WER) [53] is computed as $\text{WER}=\frac{S+D+I}{N}$, where $S$, $D$, and $I$ represent the number of substitutions, deletions, and insertions, respectively, and $N$ is the total number of words in the reference ($N=S+D+C$, with $C$ indicating correct words). A lower WER indicates better ASR performance.
@@
MDRM-test [55] Short Clips 22,208 87 short financial clip ASR WER
FinAudioSum [5] Long Audio 64 55 long financial audio Summarization ROUGE-L
@@
The evaluation dataset in FinAudio [5] is developed from two primary data sources: 1) existing open-source financial audio data originally created for non-LLM evaluation purposes. 2) novel datasets introduced in this work. Table 6 summarizes the key statistics.
  • Short financial audio clip dataset: the MDRM [55] dataset derived from earnings conference call recordings. The original audio data was segmented at the sentence level and aligned with corresponding transcripts. This dataset is initially split into training and test subsets. For our paper, we utilize only the test sets for evaluation, which includes 22,208 audio clips totaling 87 hours.
  • FinAudioSum: FinAudioSum [5] was created based on the ECTSum dataset [45], originally designed for earnings call summarization using textual data. ECTSum comprises 2,425 earnings transcripts paired with expert-generated, telegram-style summaries. We obtain corresponding audio recordings for the ECTSum test set from earningcast [^7] https://earningscast.com/. Overlapping recordings with Earnings-21 and Earnings-22 (spanning 2019–2022) are removed. The final FinAudioSum [5] dataset includes 64 recordings totaling 55 hours.
@@ Appendix H PolyFiQA-Easy and PolyFiQA-Expert
H.1 News Data Source
The data were collected from the following resources:
  • English News: All English news are collected from previous work [85].
  • Chinese News: The Chinese news were collected from public website https://guba.eastmoney.com/.
  • Japanese News: The Japanese news were mainly collected from public website https://news.mynavi.jp/ and some were generated by our own.
  • Spanish News: The Spanish news were mainly collected from public website https://www.diariolibre.com/mundo and some were generated by our own.
  • Greek News: No public available Greek news were found. All Greek news were generated by our own.
H.2 News Generation
To support high-quality, multilingual financial news generation in our benchmark, we adopt a structured semi-automated pipeline that combines human curation with large language model (LLM) assistance. The process involves three main steps:
Step 1: Extract key financial highlights
For a given company and reference date, we collect relevant financial news and corresponding SEC filings (10-Q or 10-K). From these sources, we extract 2–3 bullet points summarizing key developments. Emphasis is placed on material financial indicators such as: Revenue performance, Profit margins, and Liquidity and cash flow metrics.
Step 2: AI-assisted news drafting
We prompt GPT-4o to generate a concise financial news article using the extracted bullet points. The prompt ensures a professional and fact-based tone:

### v3

@@ E.1 English
For the IE task, it focuses on extracting structured financial information from unstructured text. We evaluate this capability using three datasets: SC Mariko et al. (2020), FinRED Sharma et al. (2022), and FINER-ORD Shah et al. (2023). The SC Mariko et al. (2020) dataset assesses the detection of causal relationships in financial news and SEC filings, using F1 Sokolova and Lapalme (2009) as the evaluation metric. The FinRED Sharma et al. (2022) dataset targets relation extraction, identifying relationships such as “product/material produced” and “manufacturer” in financial news and earnings call transcripts, also evaluated using F1 Sokolova and Lapalme (2009). The FINER-ORD Shah et al. (2023) dataset involves named entity recognition, extracting entities like organizations, locations, and persons from financial agreements and SEC filings, with performance measured by entity F1 Derczynski (2016). Overall, the IE task provides a foundation for evaluating the model’s understanding of financial knowledge.
@@
For the TA task, it encompasses tasks such as sentiment classification, topic detection, and event identification from financial texts. In this work, we focus on two representative financial datasets: Headlines Sinha and Khandait (2021) and TSA Cortis et al. (2017). The Headlines Sinha and Khandait (2021) dataset targets the extraction of actionable signals—such as price movements—from financial news headlines, with performance measured by the average F1 score (avg F1). The TSA Cortis et al. (2017) dataset centers on sentiment analysis, where the goal is to classify textual segments into positive, negative, or neutral sentiment categories, evaluated using the accuracy Makridakis (1993) metric. This task is designed to assess the capability of LLMs in performing fine-grained understanding and reasoning over domain-specific financial language.
@@
For the QA task, it involves answering specific regulatory and financial questions by leveraging the inherent knowledge of LLMs. Both FinQA Chen et al. (2021) and TATQA Zhu et al. (2021) assess models’ numerical reasoning abilities using financial reports, tables, and contexts, whereas XBRL-Math Chen et al. (2022) focus on equation inference. All are evaluated by accuracy Makridakis (1993).
@@
For the TG task, it involves summarizing or explaining specific content within regulatory or financial contexts, evaluating models’ ability to produce coherent and informative text. EDTSUM Xie et al. (2023) targets the summarization of financial news articles into concise and informative summaries. In addition, we utilize the ECTSUM Mukherjee et al. (2022) dataset, which focuses on summarizing earnings call transcripts. Model performance on these summarization tasks is evaluated using ROUGE-1 Lin (2004).
@@
For the RM task, it focuses on identifying, extracting, and analyzing risk-related information, interpreting numerical data, and evaluating a model’s ability to understand complex relationships. Specifically, we use the CCF Feng et al. (2024) datasets to assess LLMs’ ability in our paper. The CCF Feng et al. (2024) dataset is suitable for detecting fraud, which identifies whether transactions are “fraudulent” or “non-fraudulent”. This dataset is evaluated by Matthews Correlation Coefficient (MCC) Matthews (1975).
@@
For the FO task, it evaluates a model’s ability to predict future market trends and investor behavior based on patterns in financial data. We focus on stock movement prediction, aiming to forecast whether a stock’s price will rise or fall using historical price data and social media signals from the BigData22 Soun et al. (2022) dataset. Model performance is assessed using MCC Matthews (1975).
@@
For the DM task, it evaluates an LLM’s ability to synthesize diverse financial information to formulate and execute trading strategies—an inherently challenging task even for human experts. We refer to MSFT Yu et al. (2024) dataset, which consist of data over one year, simulating real-world trading scenarios using historical prices, news articles, and sentiment analysis. Model performance is assessed using Sharpe Ratio (SR) Sharpe (1998), providing a comprehensive evaluation of profitability, risk management, and strategic decision-making.
E.2 Chinese
@@
In IE task, RRE Huang et al. (2024a) is a relation extraction task that identifies the semantic relation between two audit entities in a sentence. Relations span 8 predefined categories including audit issue, audit item, audit basis, audit method, audit institution, audit outcome, audited entity, and related sector or domain. This task tests a model’s ability to infer structured relationships in financial text.
@@
For TA task, AIE Huang et al. (2024a) involves categorizing audit-target entities into 7 regulatory domains (fiscal audit, infrastructure audit, customs audit, financial audit, social security audit, tax audit, and environmental audit). This task evaluates regulatory disambiguation within Chinese financial texts. LNE Huang et al. (2024a) is a legal grounding task where models must align a given audit scenario with the most relevant legal or regulatory category (financial regulations, fiscal law, personal income tax, financial supervision, labor and employment, listed companies, social security, industry regulation, value-added tax, tax administration, asset evaluation law, resource tax, integrated governance, and comprehensive tax policy). It requires both regulatory reasoning and knowledge of Chinese legal terminology.
@@
In QA task, FinanceIQ Zhang and Yang (2023) is a multi-choice classification task where models need to answer Chinese-language financial domain-specific questions. The task evaluates an LLM’s understanding of domain-specific vocabulary, foundational financial knowledge, and contextual reasoning in Chinese.
All datasets are evaluated using classification accuracy Makridakis (1993).
E.3 Japanese
@@
To capture the unique challenges posed by Japanese in financial contexts, we only find one public TA dataset from a Japanese Financial Benchmark Hirano (2024) specifically designed to evaluate Japanese financial language understanding.
@@
For TA task, we adopt the chabsa dataset (chABSA: Aspect-Based Sentiment Analysis dataset in Japanese) Kubo et al. (2018), a binary classification task evaluates the sentiment polarity (positive or negative) of financial terms within securities reports. It is derived from publicly disclosed annual securities filings in Japan. The dataset includes annotated sentiment terms and is scored using the macro F1 Hinojosa Lee et al. (2024) metric, excluding neutral terms for evaluation consistency.
E.4 Spanish
@@
To comprehensively evaluate the understanding capabilities of LLMs in Spanish regulatory and financial contexts, we included four domain-specific datasets from FLARE-ES Zhang et al. (2024a) across three tasks: TA, QA, and TG.
@@
For TA task, text classification and sentiment analysis were included. In text classification task, the MultiFin Jørgensen et al. (2023a) dataset evaluates the model’s ability to categorize Spanish financial texts. This dataset, which centers on Spanish headlines, contains 2,066 articles spanning six key financial categories: Business & Management, Finance, Government & Controls, Industry, Tax & Accounting, and Technology. The task challenges the model to accurately assign each headline to its respective sector, thereby testing its linguistic adaptability and domain-specific knowledge in Spanish. In sentiment analysis task, the TSA Pan et al. (2023) dataset includes 3,892 entries from financial news and tweets, annotated to classify sentiments as positive, negative, or neutral—capturing a nuanced spectrum of market emotions in Spanish.
@@
For QA task, the EFPA  [^8] https://efpa-eu.org/ dataset, derived from questions used in financial examinations by official examiner associations, present unique challenges. The EFPA dataset broadens this challenge with 228 questions featuring four answer choices each, covering a broader spectrum of financial topics including economic principles, fundamental financial concepts, and intricate calculations associated with financial products. This dataset is evaluated using accuracy Makridakis (1993).
@@
For TG task, the FNS-2023 Zavitsanos et al. (2023) dataset, consisting of 232 Spanish annual reports from diverse financial companies. It aims to condense voluminous Spanish financial documents into succinct, informative abstracts, enhancing the accessibility and usability of financial information. Rich in detailed financial data and narratives, these reports pose a distinct challenge: distilling the essence of each document into a summary that captures the key information while preserving the factual integrity and coherence of the original text. This dataset is evaluated using ROUGE-1 Lin (2004).
E.5 Greek
@@
We adopt four core financial NLP datasets from Plutus-ben Peng et al. (2025), the first Greek financial benchmark. These datasets span from IE, TA, QA, and TG tasks.
@@
For IE task, only GRFinNUM Peng et al. (2025) is included. GRFinNUM Peng et al. (2025) targets fine-grained classification of numerals into MONETARY, PERCENTAGE, TEMPORAL, QUANTITY, and OTHERS categories, which captures the nuances of numerical semantics in long-form Greek financial texts. This dataset is evaluated using entity F1 Derczynski (2016).
@@
For TA task, GRMultiFin Jørgensen et al. (2023b) focuses on classifying concise Greek financial headlines into one of six predefined thematic categories: Business & Management, Finance, Government & Controls, Tax & Accounting, Technology, and Industry, emphasizing the need for lexical disambiguation and contextual inference due to the brevity and ambiguity typical of headlines. This dataset is evaluated using accuracy Makridakis (1993).
@@
For QA task, GRFinQA Peng et al. (2025) evaluates models’ performance on inferring the correct answer using provided financial text under a multiple-choice format, assessing a model’s ability to comprehend and reason over finance-related questions drawn from Greek academic sources. This dataset is evaluated using accuracy Makridakis (1993).
@@
For TG task, GRFNS-2023 Zavitsanos et al. (2023) tests models’ ability of generating fluent, coherent, and factually accurate summaries from complex narrative sections of Greek financial reports, testing a model’s capacity for paraphrasing and information compression. This dataset is evaluated using ROUGE-1 Lin (2004).
Appendix F Audio Benchmark
@@
The Automatic Speech Recognition (ASR) task evaluates AudioLLMs’ accuracy in transcribing financial audio clips, directly impacting applications such as financial voice assistants. In this task, each input is represented as $(A,Q,R)$, where $A$ is an audio clip, $Q$ is the prompt instruction, and $R$ is the reference transcript. The transcribed text $T$ is generated as $T=\textit{AudioLLM}(A,Q)$. The Word Error Rate (WER) Park et al. (2008) is computed as $\text{WER}=\frac{S+D+I}{N}$, where $S$, $D$, and $I$ represent the number of substitutions, deletions, and insertions, respectively, and $N$ is the total number of words in the reference ($N=S+D+C$, with $C$ indicating correct words). A lower WER indicates better ASR performance.
@@
MDRM-test Qin and Yang (2019) Short Clips 22,208 87 short financial clip ASR WER
FinAudioSum Cao et al. (2025b) Long Audio 64 55 long financial audio Summarization ROUGE-L
@@
The evaluation dataset in FinAudio Cao et al. (2025b) is developed from two primary data sources: 1) existing open-source financial audio data originally created for non-LLM evaluation purposes. 2) novel datasets introduced in this work. Table 6 summarizes the key statistics.
  • Short financial audio clip dataset: the MDRM Qin and Yang (2019) dataset derived from earnings conference call recordings. The original audio data was segmented at the sentence level and aligned with corresponding transcripts. This dataset is initially split into training and test subsets. For our paper, we utilize only the test sets for evaluation, which includes 22,208 audio clips totaling 87 hours.
  • FinAudioSum: FinAudioSum Cao et al. (2025b) was created based on the ECTSum dataset Mukherjee et al. (2022), originally designed for earnings call summarization using textual data. ECTSum comprises 2,425 earnings transcripts paired with expert-generated, telegram-style summaries. We obtain corresponding audio recordings for the ECTSum test set from earningcast [^9] https://earningscast.com/. Overlapping recordings with Earnings-21 and Earnings-22 (spanning 2019–2022) are removed. The final FinAudioSum Cao et al. (2025b) dataset includes 64 recordings totaling 55 hours.
Appendix G OCR Statistics
Dataset # PDF Files Storage (GB) Size
EnglishOCR 1,039 1.70 7,961
JapaneseOCR 331 1.57 17,586
SpanishOCR 2,302 0.76 12,819
GreekOCR 100 0.16 6,533
Table 7: Statistics of OCR datasets.
@@ Appendix H PolyFiQA-Easy and PolyFiQA-Expert
H.1 Data Construction
Figure 8: Expert-in-the-loop data construction of PolyFiQA-Easy and PolyFiQA-Expert.
Expert-in-the-Loop Data Construction.
To ensure benchmark fidelity and domain rigor, we adopt an expert-in-the-loop pipeline (Figure 8). Three financial professionals (Appendix H.7) with expertise in economics, business, and accounting oversaw all phases of construction, including news selection, question authoring, guideline development, annotation, and quality control. News articles were meticulously screened for strong alignment with financial reports. Questions were crafted to anchor in real analytical tasks and span two difficulty tiers: easier questions in PolyFiQA-Easy and more complex ones in PolyFiQA-Expert (Appendix H.4). This structure supports fine-grained model assessment across reasoning levels (Appendix H.5). Rigorous tier-specific annotation guidelines (Appendix H.6) were refined through iterative pilot annotation rounds including tier-specific annotation rules and formatting protocols to promote inter-annotator consistency. In total, approximately 130 expert hours were logged in Label Studio (Appendix H.8), establishing a high quality, reproducible, auditable, and streamlined workflow aligned with best practices in benchmark creation. The raw datasets were further converted into structured instruction datasets with task-specific prompts thoughtfully crafted by financial professionals (Appendix H.11).
Quality Validation.
Evaluating the quality of free-text generation datasets remains a persistent challenge in both QA and TG. To ensure annotation reliability, we adopt a structured scoring framework inspired by prior summarization benchmarks Takeshita et al. (2024); Gliwa et al. (2019), evaluating each instance along two key dimensions: Relevance (scored 1–4) captures whether the response includes the key information required to answer the question. Consistency (scored 1–3) measures factual accuracy, especially numerical values. Each question is initially annotated by one expert and independently scored by two additional reviewers using detailed, pilot-refined validation guidelines (Appendix H.9). Only responses with cumulative scores above 5 are retained in the final dataset. To further validate scoring reliability, we report inter-annotator agreement as normalized difference percentage across dimensions. PolyFiQA-Easy and PolyFiQA-Expert achieved average inter-annotator agreements of 89.38% and 91.21%, demonstrating the benchmark’s high quality and scoring consistency.
H.2 News Data Sources
We collected news data from the following resources:
H.3 News Generation
To support high-quality, multilingual financial news creation in our benchmark, we designed a semi-automated pipeline that integrates human curation with LLM assistance. The pipeline consists of three main stages:
Step 1: Extracting key financial highlights.
For each target company and reference date, we collect relevant financial news and the corresponding SEC filings (10-Q or 10-K). From these sources, we distill 2–3 bullet points summarizing key developments, with emphasis on material indicators such as revenue performance, profit margins, and liquidity or cash flow metrics.
Step 2: AI-assisted drafting.
We prompt GPT-4o to generate a concise financial news article based on the extracted bullet points. The prompt enforces a neutral and professional journalistic tone, while discouraging unsupported statements: